### v1

# The $B^{+}\to K^{+}\nu\bar{\nu}$ decay as a search for the QCD axion

Abstract
We reinterpret the $B^{+}\to K^{+}\nu\bar{\nu}$ measurement as a probe of the decay $B^{+}\to K^{+}a$, where $a$ denotes an axion or axion-like particle. This is possible in the kinematic regions where the di-neutrino invariant mass squared, $q^{2}$, can be identified with the assumed $m_{a}^{2}$. While $q^{2}$ is reconstructed only in the Hadronic Tag Analysis (HTA), it is not directly accessible in the Inclusive Tag Analysis (ITA), which provides much larger statistics and hence higher sensitivity. In ITA, $q^{2}$ is proxied by a reconstructed variable, $q^{2}_{\text{rec}}$, usually modeled through Belle II simulations. We show that the mapping between $q^{2}$ and $q^{2}_{\text{rec}}$ can be accurately derived from kinematic arguments alone—without relying on internal experimental inputs. Using this relation and publicly available efficiencies, we obtain the strongest existing bounds on the coupling-rescaled Peccei-Quinn scale, $|(F_{V})_{sb}|\geq 0.9\times 10^{9}$ GeV, improving the latest bound in the literature by an order of magnitude. We further show that the bound depends only marginally on the assumed value of the di-neutrino signal strength $\mu_{\nu\bar{\nu}}$, whether fixed or floated. This establishes $B^{+}\to K^{+}\nu\bar{\nu}$ as a double probe—of new short-distance physics in the $B^{+}\to K^{+}\nu\bar{\nu}$ amplitude, and of new light, elusive particles produced via $B^{+}\to K^{+}a$—the two probes working independently to an excellent approximation.
preprint: LAPTH-036/25
Introduction— Spin-zero, odd-parity particles below the GeV scale—collectively known as axions or axion-like particles (ALPs)—are hypothetical states actively sought through laboratory experiments [1, 2, 3], astrophysical data [4, 5, 6], cosmological probes [7], and theoretical studies [8, 9, 10]. A central focus of this research domain is the “invisible” [11, 12, 13, 14] QCD axion [15, 16, 17, 18], henceforth the axion, which elegantly solves the strong-$CP$ problem—the unexplained suppression of a $CP$-violating QCD term. Given the extreme suppression, $O(10^{-10})$, neither coincidence nor anthropic reasoning provides a satisfactory explanation [19, 20]. The axion also offers a viable Dark Matter candidate [21, 22, 23]. Unlike generic ALPs, its mass $m_{a}$ and decay constant $f_{a}$ are linked: smaller $m_{a}$ implies weaker couplings, scaling as $1/f_{a}$.
Existing bounds indicate that $m_{a}$ must be several orders of magnitude below the pion mass [9]. Yet this small mass also opens the possibility that axions could be copiously produced at high-intensity colliders, without having to resort to high energies. One can consider decays of $K$, $D$, and $B$ mesons, or even the baryonic counterparts, decaying to fully reconstructible final states, plus the axion, which is assumed to escape without decay nor interaction within the detector volume. This represents the most general and conservative scenario.
Since the QCD axion manifests as an energy and momentum imbalance—equal in magnitude, as the axion has negligible mass with respect to the lightest physical scale of the decay—it is crucial that the remaining final-state particles be reconstructible with high precision. Moreover, the identity and momentum of the decaying hadron must be accurately known to better control the many backgrounds involving elusive particles that can mimic the axion.
Belle II [24] is particularly well positioned to satisfy both of these constraints. It collides $e^{-}$ and $e^{+}$ beams with momenta in approximately opposite directions and magnitudes of $7$ and $4$ GeV  [^1] The spread, or bite, around these momenta is tiny and completely negligible for the purposes of this paper., respectively. The center-of-mass energy corresponds to the $b\bar{b}$ resonance known as $\Upsilon(4S)$, which subsequently decays with near-unit probability into charged or neutral $B\bar{B}$ meson pairs. Signals are then studied in the decays of one of the produced mesons, $B_{\mathrm{s}ig}$, while the other meson is often used as a tag decay—allowing the kinematics of the signal decay to be fully reconstructed and the flavour of $B_{\mathrm{s}ig}$ to be identified.
In this paper we will consider the $B^{+}\to K^{+}a$ decay  [^2] The conjugated mode with negatively-charged mesons is assumed to be included throughout our study.. The axion-hadron interaction responsible for the decay inherits in a calculable way from the underlying axion-quark interaction, which just below the electroweak scale can be written most generally as [27]
$\mathcal{L}_{a}\supset\;$ $\frac{\partial_{\mu}a}{2f_{a}}\sum_{i,j=1}^{3}\bar{d}_{i}\gamma^{\mu}\bigl[(k_{V})_{d_{i}d_{j}}+\gamma_{5}(k_{A})_{d_{i}d_{j}}\bigl]d_{j}+\left\{d\to u\right\}$
$-\frac{\alpha_{s}}{8\pi}\frac{a}{f_{a}}G\tilde{G}~,$ (1)
where $u_{i},d_{i}$ denote up- and down-type quarks of generation $i$ in the mass eigenbasis. The $G\tilde{G}$ operator  [^3] The sign difference in the $f_{a}$-defining $aG\tilde{G}$ term in Eq.The $B^{+}\to K^{+}\nu\bar{\nu}$ decay as a search for the QCD axion with respect to Ref.[9] arises from our use of opposite $\epsilon$-tensor conventions. can be traded for an axion-field dependent quark mass matrix by redefining the quark fields through $q\to\exp[i(Q_{V}+Q_{A}\gamma_{5})\,a/(2f_{a})]q$, with $Tr Q_{A}=1$ [27]. Although the hermitian matrices $Q_{V},Q_{A}$ can be chosen as diagonal, they introduce an unphysical reparametrization ambiguity [27].
The process $B^{+}\to K^{+}a$, on which this paper is focused, involves the complex coupling $(k_{V})_{sb}$. The calculation of this process requires the following hadronic matrix element
$\langle K^{-}(k)|\bar{s}b|B^{-}(p)\rangle=\frac{m_{B}^{2}-m_{K}^{2}}{m_{b}-m_{s}}f_{0}(m_{a}^{2})~.$ (2)
Here, form-factor values and conventions are taken from Ref. [29]. The relation in Eq. (2) imply the following branching ratios (BRs) [30, 31]
$\mathcal{B}(B^{-}\to K^{-}a)=\frac{\tau_{B}|(k_{V})_{sb}|^{2}\,|\vec{k}|m_{B}^{2}}{32\pi f_{a}^{2}}\left(1-\frac{m_{K}^{2}}{m_{B}^{2}}\right)^{2}f_{0}^{2}(m_{a}^{2})~,$ (3)
where $|\vec{k}|=\lambda^{1/2}(m_{B}^{2},m_{K^{(*)}}^{2},m_{a}^{2})/(2m_{B})$ and $\lambda(a,b,c)=a^{2}+b^{2}+c^{2}-2ab-2ac-2bc$. While the theoretically well-motivated QCD axion requires $m_{a}\simeq 0$, we will also discuss the general case of an ALP. Here and henceforth, the axion-quark couplings are understood at a scale close to $m_{B}$.
We search for the signal Eq. (3) in data collected by Belle II for the $B^{+}\to K^{+}\nu\bar{\nu}$ measurements. Belle II employs two nearly uncorrelated datasets, analyzed using the “classic” Hadronic Tagging Analysis (HTA) method and the more recent Inclusive Tagging Analysis (ITA) method, respectively [32, 33, 34]. The ITA method achieves much larger statistics and thus dominates the sensitivity to the measurement. Reliably recasting this dataset for our search thus hinges on one key piece of non-public information: the mapping between the true di-neutrino invariant mass squared, $q^{2}$, which is inaccessible within the ITA method, and its reconstructed counterpart, $q^{2}_{{\mathrm{r}ec}}$, defined below. We show that the $q^{2}$ distribution as a function of $q^{2}_{{\mathrm{r}ec}}$ can be reliably inferred from kinematic arguments alone—i.e., without requiring internal information from the experiment. Using this mapping and publicly available efficiencies, we demonstrate that this search yields the strongest existing bounds on the magnitudes of the fundamental QCD-axion coupling $(k_{V})_{sb}$.
Both of the aforementioned ITA and HTA methods yield a central value in excess of the SM prediction [35, 36]. This excess motivates the following parameterization of the BR for $B^{+}\to K^{+}\nu\bar{\nu}$—differential in the squared di-neutrino invariant mass $q^{2}$
$d\mathcal{B}_{\nu\bar{\nu}}(\mu)/dq^{2}~\equiv~\mu_{\nu\bar{\nu}}\,d\mathcal{B}_{\nu\bar{\nu},{\mathrm{S}M}}/dq^{2}~,$ (4)
where on the r.h.s. we use the SM prediction [35, 37, 38], rescaled by a signal-strength factor. This rescaling allows one to disentangle the presence of new effects in the $B^{+}\to K^{+}\nu\bar{\nu}$ mode from those due to the two-body decay $B^{+}\to K^{+}a$, and to test whether the bound inferred on the $(k_{V})_{sb}$ coupling is affected by possible new-physics contamination in the $B^{+}\to K^{+}\nu\bar{\nu}$ sample.
Within the ITA, Belle II cannot directly reconstruct $q^{2}\equiv m_{\nu\bar{\nu}}^{2}=(p-k)^{2}$ and thus constructs a proxy variable, $q^{2}_{{\mathrm{r}ec}}$ [32], in terms of the known $s\equiv(p_{e^{-}}+p_{e^{+}})^{2}$ and the energy of the final-state kaon in the $B\bar{B}$ rest frame (hereafter denoted with a star). In this frame, the signal $B$-meson momentum is known in magnitude, $|\vec{p}_{B}^{\,*}|=0.3298~{\mathrm{G}eV}$, and is distributed isotropically. In turn, in the $B$ rest frame (denoted with a double star), the kaon is emitted isotropically. One can therefore write
$\begin{split}q^{2}_{{\mathrm{r}ec}}~\equiv~&\frac{s}{4}+m_{K}^{2}-\sqrt{s}E_{K}^{*}~=~q^{2}+E_{B}^{*2}-m_{B}^{2}-2\,\vec{p}_{B}^{\,*}\cdot\vec{p}_{K}^{\,*}(q^{2})\\ ~=~&q^{2}+E_{B}^{*2}-m_{B}^{2}-2|\vec{p}_{B}^{\,*}|\gamma^{**}\left(\beta^{**}E_{K}^{**}(q^{2})+|\vec{p}_{K}^{\,**}(q^{2})|\cos\theta^{**}\right)~,\end{split}$ (5)
where $|\vec{p}_{K}^{\,**}|=\lambda^{1/2}(m_{B}^{2},m_{K}^{2},q^{2})/(2m_{B})$ and $\beta^{**}=|\vec{p}_{B}^{\,*}|/E_{B}^{*}=0.06234$ is the boost parameter in the Lorentz transformation from the $B$ rest frame to the $B\bar{B}$ rest frame. Finally, $\theta^{**}$ is the polar angle of the kaon momentum in the $B$-meson rest frame relative to the $B$-meson momentum in the $B\bar{B}$ rest frame. Importantly, $\theta^{**}$ constitutes the only unmeasured quantity in Eq. (5). The distribution of $q^{2}_{{\mathrm{r}ec}}$ for a given $q^{2}$ can therefore be derived from Eq. (5) with $c^{**}\equiv\cos\theta^{**}\in[-1,1]$ uniformly distributed. The sub-percent-level uncertainties [39] smearing the energies and momenta that appear in Eq. (5) have a completely negligible impact on the resulting $q^{2}(q^{2}_{{\mathrm{r}ec}})$ distribution.
The physical branching ratios and the observed yields are then related in the ITA as
$\frac{dN_{X}}{dq^{2}_{{\mathrm{r}ec}}}~=~N_{B}\int dq^{2}\int dc^{**}\,\frac{1}{2}\,\delta\Bigl(q^{2}_{{\mathrm{r}ec}}-q^{2}_{{\mathrm{r}ec}}(q^{2},c^{**})\Bigl)\,\epsilon_{\text{ITA}}(q^{2})\,\frac{d\mathcal{B}_{X}}{dq^{2}}~\equiv~N_{B}\int dq^{2}\,f_{q^{2}_{{\mathrm{r}ec}}}(q^{2})\,\epsilon_{\text{ITA}}(q^{2})\,\frac{d\mathcal{B}_{X}}{dq^{2}}~,$ (6)
where $N_{B}=387(6)\times 10^{6}$ [32]  [^4] The below-2% $N_{B}$ error has a negligible impact on our analysis, because $n_{\nu\bar{\nu},{\mathrm{b}kg}}^{i}$ have an error approaching 10%. is the number of $B$ candidates in the given decay channel $X$, $\mathcal{B}$ the corresponding BR, and $\epsilon_{\text{ITA}}(q^{2})$ the ITA-specific selection efficiency as a function of $q^{2}$, which can be obtained from Ref. [32]. The function $q^{2}_{{\mathrm{r}ec}}(q^{2},c^{**})$ appearing in the delta function is the r.h.s. of Eq. (5). In the last equality we make contact with the smearing function $f_{q^{2}_{{\mathrm{r}ec}}}$ [41, 31, 42], which is typically assumed to require Belle II simulation.
The $q^{2}$ vs. $q^{2}_{{\mathrm{r}ec}}$ correlation implied by Eq. (6) is shown in Fig. 1.
Figure 1: Correlation between the true ($q^{2}$) and the reconstructed ($q^{2}_{{\mathrm{r}ec}}$) di-neutrino invariant mass squared. The plot is normalised to $10^{5}$ observed events.
This correlation plot may be compared with figure 5.1 of Ref. [43], which is the output of internal Belle II simulations. The two figures display good agreement, aside from an inconsequential low-$q^{2}$ tail visible in figure 5.1 of Ref. [43]  [^5] This tail originates from radiative decays with undetected photons. Strictly speaking, $B\to K\nu\bar{\nu}(+n\gamma)$ is a distinct process from $B\to K\nu\bar{\nu}$ and should be subtracted if sizeable, as in $B_{s}\to\mu^{+}\mu^{-}$[50]. Here, however, the effect is negligible (well below 1% in Ref. [43]), since the lightest emitter is the $K^{+}$, which makes the exponent $b$ in ${\mathcal{B}}(B\to K\nu\bar{\nu}(+n\gamma))=\Bigl(\tfrac{2\Delta E}{m_{B}}\Bigr)^{\tfrac{\alpha}{4\pi}b}{\mathcal{B}}(B\to K\nu\bar{\nu})$ parametrically small. Here $\Delta E$ denotes the minimal photon energy in the $B$ rest frame required for detectability—a “theoretical” scale usually proxied by the $\sqrt{q^{2}}$ resolution. .
To make the origin of the $q^{2}_{{\mathrm{r}ec}}$ smearing explicit, we recall that the relation between $q^{2}$ and $q^{2}_{{\mathrm{r}ec}}$, Eq. (5), introduces a uniformly distributed variable from an unmeasured angle. This kinematic unknown constitutes an irreducible source of uncertainty, which is not immediately transparent in simulation output. The good agreement between Fig. 1 and figure 5.1 of Ref. [43] shows that this mechanism alone accounts for the observed correlation, with any additional effects being negligible.
Within the HTA, instead, Belle II can directly reconstruct the $q^{2}$ of the events, so the physical BRs and observed yields are simply related as
$\frac{dN_{X}}{dq^{2}}~=~N_{B}\,\epsilon_{\text{HTA}}(q^{2})\,\frac{d\mathcal{B}_{X}}{dq^{2}}~,$ (7)
where $\epsilon_{\text{HTA}}(q^{2})$ is the HTA-specific selection efficiency as a function of $q^{2}$, also available in Ref. [32].
Analysis— To probe the $B^{+}\to K^{+}a$ decay through the $B^{+}\to K^{+}\nu\bar{\nu}$ measurements at Belle II, we construct a likelihood ratio test statistic based solely on the public data of Fig. 17 (ITA) and Fig. 20 (HTA) of Ref. [32]. The likelihood function $\mathcal{L}$ is constructed as the product of two contributions, i.e. ${\mathcal{L}}={\mathcal{L}}_{1}{\mathcal{L}}_{2}$.
The first and main term in the likelihood function represents a Poisson distribution $P$ for each bin, labeled by an index $i$, assuming the bins to be uncorrelated: ${\mathcal{L}}_{1}=\prod_{i}^{n_{\mathrm{b}in}}P(n^{i}_{\mathrm{o}bs};\lambda^{i})$. Here $n^{i}_{\mathrm{o}bs}$ is the number of observed events in bin $i$, defined by $q^{2}\in[q^{2}_{i},q^{2}_{i+1})$, and $\lambda^{i}$ is the corresponding expected number of events, which depends on $\mathcal{B}(B^{+}\to K^{+}a)\equiv\mathcal{B}_{a}$ and on the signal strength $\mu_{\nu\bar{\nu}}$—the two key quantities to be determined:
$\begin{split}\lambda^{i}(\mathcal{B}_{a},\mu_{\nu\bar{\nu}})~=~\mathcal{B}_{a}\,n^{i}_{a}+\mu_{\nu\bar{\nu}}\,n^{i}_{\nu\bar{\nu}}+(1+\sigma^{\text{bkg}}\,\theta)\,n^{i}_{\text{bkg}}~.\end{split}$ (8)
This expression applies to either the ITA or HTA dataset, where $n^{i}_{a}$, $n^{i}_{\nu\bar{\nu}}$, and $n^{i}_{\mathrm{b}kg}$ are the yields in the $i^{\mathrm{t}h}$ bin for the signal, the $B^{+}\to K^{+}\nu\bar{\nu}$ decay, and the background, respectively. Within the two methods, the first two components on the r.h.s. of Eq. (8) are defined as:
$n^{i}_{a,\text{ITA}}=$ $N_{B}\int\limits_{\text{$i^{\text{th}}$ bin}}dq^{2}_{{\mathrm{r}ec}}\,f_{q^{2}_{{\mathrm{r}ec}}}(m_{a}^{2})\,\epsilon_{\text{ITA}}(m_{a}^{2})~,$ (9)
$n^{i}_{\nu\bar{\nu},\text{ITA}}=$ $N_{B}\int\limits_{\text{$i^{\text{th}}$ bin}}dq^{2}_{{\mathrm{r}ec}}\int dq^{2}\,f_{q^{2}_{{\mathrm{r}ec}}}(q^{2})\,\epsilon_{\text{ITA}}(q^{2})\,\frac{d\mathcal{B}_{\nu\bar{\nu},{\mathrm{S}M}}}{dq^{2}}~,$ (10)
$n^{i}_{a,\text{HTA}}=$ $N_{B}\int\limits_{\text{$i^{\text{th}}$ bin}}dq^{2}\,\epsilon_{\text{HTA}}(q^{2})\,g_{a}(q^{2})~,$ (11)
$n^{i}_{\nu\bar{\nu},\text{HTA}}=$ $N_{B}\int\limits_{\text{$i^{\text{th}}$ bin}}dq^{2}\,\epsilon_{\text{HTA}}(q^{2})\,\frac{d\mathcal{B}_{\nu\bar{\nu},{\mathrm{S}M}}}{dq^{2}}~.$ (12)
These expressions can be understood in terms of Eqs. (6)-(7), bearing in mind that for the signal case the $q^{2}$ dependence of the differential BR is a delta function with support at $q^{2}=m_{a}^{2}$, and the overall normalization $\mathcal{B}_{a}$ is factored out in Eq. (8). In the HTA case, one has to take into account that the invariant mass resolution of the experiment is finite. We do so by replacing $\delta(q^{2}-m_{a}^{2})$ with a normal distribution $g_{a}(q^{2})$ centered at the axion mass value with uncertainty taken to be $\sim(450~{\mathrm{M}eV})^{2}$, comparable to the beam-constrained mass resolution of $B$ mesons reconstructed using the hadronic tag [45]  [^6] In principle, the same smearing should also be applied to the di-neutrino channel, Eq. (12). In the axion case this treatment is essential, as it broadens the Dirac-delta shape of the $q^{2}$ distribution. For the di-neutrino mode, however, the impact is much milder because its $q^{2}$ spectrum is already smooth. Such an effect would become relevant only once a measurement of the differential $\mathcal{B}_{\nu\bar{\nu}}$ reaches comparable precision. In practice, the smeared and unsmeared $q^{2}$ distributions differ only marginally: a few events with small positive $q^{2}$ migrate to slightly negative reconstructed values, thus shifting between adjacent bins of the Belle II spectrum. Numerically, the neutrino likelihood at $\mathcal{B}_{a}=0$ remains nearly unchanged, and the bound on the axion is identical within the two-digit precision considered..
As regards the last component of Eq. (8), the numbers of expected background events for each bin, $n^{i}_{\mathrm{b}kg}$, can be read out from the plots of the collaboration. While the uncertainty on the Montecarlo generated background events are not explicitly provided, we infer that the uncertainty on the total normalisation is respectively $\sigma_{\text{ITA}}^{\text{bkg}}\approx 1\%$ and $\sigma_{\text{HTA}}^{\text{bkg}}\approx 10\%$. These estimates have been obtained from the $q^{2}$ distribution and its corresponding pull shown in Figs. 17 and 20 of Ref. [32]. One can in fact identify the pull with $|n^{i}_{\mathrm{d}at}-n^{i}_{\mathrm{h}st}|/(\sqrt{(\sigma^{i}_{\mathrm{d}at})^{2}+(\sigma^{i}_{\mathrm{h}st})^{2}})$, where the subscripts ‘dat’ and ‘hst’ denote respectively the data and the value of the stacked histograms, and $\sigma^{i}_{\mathrm{d}at}$ the uncertainty in the data point. Inverting the pull distribution in terms of $\sigma^{i}_{\mathrm{h}st}$ provides an estimate of the above uncertainty. Thus, we introduced two nuisance parameters, i.e. $\theta_{\text{ITA}}$ and $\theta_{\text{HTA}}$, both constrained by a normal Gaussian distribution in the likelihood as
${\mathcal{L}}_{2}=\frac{1}{\sqrt{2\pi}}e^{-\theta^{2}_{\text{ITA}}/2}\times\frac{1}{\sqrt{2\pi}}e^{-\theta^{2}_{\text{HTA}}/2}\,.$ (13)
As a non-trivial check of our procedure, we derive the profile likelihood ratio as a function of the signal strength $\mu_{\nu\bar{\nu}}$ in the case the axion is absent, i.e. ${\mathcal{B}}_{a}=0$, obtaining very consistent results as compared with the collaboration’s counterparts. Details are relegated to the text around Fig. 3 in the Supplemental Sec. A.
[…]
ITA HTA Combined
$\mu_{\nu\bar{\nu}}=1$ $n_{\text{dof}}=2$ $n_{\text{dof}}=2$ $n_{\text{dof}}=3$
$\mathcal{B}_{a}<1.1\times 10^{-6}$ $\mathcal{B}_{a}<5.1\times 10^{-6}$ $\mathcal{B}_{a}<1.4\times 10^{-6}$
$|(F_{V})_{sb}|>8.4\times 10^{8}$ GeV $|(F_{V})_{sb}|>3.9\times 10^{8}$ GeV $|(F_{V})_{sb}|>7.6\times 10^{8}$ GeV
$\mu_{\nu\bar{\nu}}\neq 1$ $n_{\text{dof}}=3$ $n_{\text{dof}}=3$ $n_{\text{dof}}=4$
$\mathcal{B}_{a}<1.0\times 10^{-6}$ $\mathcal{B}_{a}<6.2\times 10^{-6}$ $\mathcal{B}_{a}<1.2\times 10^{-6}$
$|(F_{V})_{sb}|>8.9\times 10^{8}$ GeV $|(F_{V})_{sb}|>3.6\times 10^{8}$ GeV $|(F_{V})_{sb}|>8.0\times 10^{8}$ GeV
Table 1: Upper limits on the axion BR and lower limits on the coupling-rescaled Peccei-Quinn scale at 90% C.L. for the different cases considered. All limits assume $m_{a}=0$.
The different cases described and the resulting bounds are summarized in Table 1, which assumes $m_{a}=0$. Results for a general value of the axion mass are shown in Fig. 2. This figure quantitatively conveys two important pieces of information. First, as one would expect, the bound is weakened by an order of magnitude in the $m_{\nu\bar{\nu}}$ bins where data and the SM $B^{+}\to K^{+}\nu\bar{\nu}$ prediction are in tension. These discrepant bins can actually be used to set both an upper and a lower bound on $\mathcal{B}_{a}$ [47, 31, 42], as shown in Fig. 4 in the Supplemental Sec. A.
Second, and less obviously, the bound depends only marginally on the assumption of new physics in $\mathcal{B}_{\nu\bar{\nu}}$—namely, on whether $\mu_{\nu\bar{\nu}}$ is set to unity or left floating. This is an important conclusion, making the $B^{+}\to K^{+}\nu\bar{\nu}$ decay a double probe: of new short-distance physics in the $B^{+}\to K^{+}\nu\bar{\nu}$ amplitude, testable by $\mu_{\nu\bar{\nu}}\neq 1$, and of new light, elusive particles, produced as $B\to Ka$, with a BR sensitivity proportional to the total $\mathcal{B}_{\nu\bar{\nu}}$ uncertainty [48]. This second point is also conveyed by the similarity between the bounds in the first and second rows of Table 1. In the case where $\mathcal{B}_{\nu\bar{\nu}}$ is fixed to its SM value, we verify that varying it by as much as 10% has a negligible impact on these results. To corroborate this conclusion even further, we evaluate the bound on $\mathcal{B}_{a}$ assuming fixed values of $\mu_{\nu\bar{\nu}}$ but different from unity. As shown in Fig. 5 of the Supplemental Sec. A, the resulting bound has only a mild dependence on the specific value of the di-neutrino signal strength—even for huge beyond-SM contributions to the $B^{+}\to K^{+}\nu\bar{\nu}$ amplitude.
Figure 2: Upper limit on $\mathcal{B}_{a}$ (upper row) and lower limit on the coupling-rescaled Peccei-Quinn scale $(F_{V})_{sb}$ (lower row) at 90% CL as a function of the mass $m_{a}$, and with $\mu_{\nu\bar{\nu}}$ set to unity (blue line) or left floating (red).
The BR limit in Fig. 2 (upper row) can be translated into a bound on the coupling-rescaled Peccei-Quinn scale $(F_{V})_{sb}\equiv 2f_{a}/(k_{V})_{sb}$ [49], to obtain a bound on this scale. This bound is shown in Fig. 2 (lower row), as a function of $m_{a}$ for the two alternative $\mu_{\nu\bar{\nu}}$ possibilities. Limits for $m_{a}\simeq 0$ are collected in Table 1. As the table shows, we obtain $|(F_{V})_{sb}|\geq 0.9\times 10^{9}$ GeV for the QCD axion, which improves by one order of magnitude the pre-existing bound [49].
In conclusion, we reinterpret the $B^{+}\to K^{+}\nu\bar{\nu}$ measurement as a probe of $B^{+}\to K^{+}a$, with $a$ an axion ($m_{a}\simeq 0$) or ALP ($m_{a}>0$). We derive analytically the mapping between the measured $q^{2}_{\text{rec}}$ and the true $q^{2}$, allowing the Inclusive Tag Analysis to constrain axion emission without relying on internal simulations. Using public efficiencies, we set the strongest existing limit on the Peccei–Quinn scale rescaled by the fundamental axion coupling relevant to our channel. The result is largely independent of the assumed $\mu_{\nu\bar{\nu}}$, making $B^{+}\to K^{+}\nu\bar{\nu}$ a dual probe of short-distance new physics in the $B^{+}\to K^{+}\nu\bar{\nu}$ amplitude and of the production of an undetected axion. This approach can be generalized to $B^{+}\to K^{*+}\nu\bar{\nu}$—and to yet other channels—as soon as Belle II measurements become available.
Acknowledgments— We thank Joel C. Swallow for useful discussions. This work has received funding from the French ANR, under contracts ANR-19-CE31-0016 (‘GammaRare’) and ANR-23-CE31-0018 (‘InvISYble’), that we gratefully acknowledge.
A Supplementary plots
In this part we collect additional plots which are not necessary to understand the results in the main text, but further corroborate them, or can be used to perform comparisons.
Fig. 3 shows the log-likelihood profiles resulting from our analyses and their comparison with the Belle II counterparts. As the figure indicates, we obtain results very close to those of the collaboration, except in the region $\mu_{\nu\bar{\nu}}<1$ for the HTA case. This small deviation is understood: it originates from our treatment of the background uncertainty. When the background is fixed to its central value with no associated error, the minimum coincides with the Belle II result; including the background error, however, broadens the minimum and slightly lowers the best-fit point.
Fig. 4 shows $\mathcal{B}_{a}$ as a signal, displaying the $1\sigma$, $2\sigma$, and $3\sigma$ regions for $\mathcal{B}_{a}$ as a function of the mass, with $\mu_{\nu\bar{\nu}}$ either fixed to unity or left floating.
In Fig. 5, we evaluate the bound on $\mathcal{B}_{a}$ for fixed values of $\mu_{\nu\bar{\nu}}$ different from unity, illustrating the mild dependence of the resulting bound on the assumed value of the di-neutrino signal strength.
Figure 3: Twice the negative profile log-likelihood ratio as a function of $\mu_{\nu\bar{\nu}}$ for the ITA (red), HTA (blue) or combined (black) analyses. The solid line refers to our results while the dashed line are the results from Belle II in Fig. 16 of Ref. [32]. The grey dashed lines refer to $1,2,3\sigma$ $\chi^{2}$ values with 1 degree of freedom.
Figure 4: Regions $\mathcal{B}_{a}$ at 68% (green), 95% (yellow) and 99% (orange) CL and best fit value (solid green line) of $\mathcal{B}_{a}$ as a function of the mass with $\mu_{\nu\bar{\nu}}$ fixed to 1 (upper row) or left floating (lower row).
Figure 5: Upper limit at 90% CL on $\mathcal{B}_{a}$ for fixed values of $\mu_{\nu\bar{\nu}}$.

### v2

@@
 Abstract
-We reinterpret the $B^{+}\to K^{+}\nu\bar{\nu}$ measurement as a probe of the decay $B^{+}\to K^{+}a$, where $a$ denotes an axion or axion-like particle.
-This is possible in the kinematic regions where the di-neutrino invariant mass squared, $q^{2}$, can be identified with the assumed $m_{a}^{2}$.
-While $q^{2}$ is reconstructed only in the Hadronic Tag Analysis (HTA), it is not directly accessible in the Inclusive Tag Analysis (ITA), which provides much larger statistics and hence higher sensitivity.
-In ITA, $q^{2}$ is proxied by a reconstructed variable, $q^{2}_{\text{rec}}$, usually modeled through Belle II simulations.
-We show that the mapping between $q^{2}$ and $q^{2}_{\text{rec}}$ can be accurately derived from kinematic arguments alone—without relying on internal experimental inputs.
-Using this relation and publicly available efficiencies, we obtain the strongest existing bounds on the coupling-rescaled Peccei-Quinn scale, $|(F_{V})_{sb}|\geq 0.9\times 10^{9}$ GeV, improving the latest bound in the literature by an order of magnitude.
-We further show that the bound depends only marginally on the assumed value of the di-neutrino signal strength $\mu_{\nu\bar{\nu}}$, whether fixed or floated.
-This establishes $B^{+}\to K^{+}\nu\bar{\nu}$ as a double probe—of new short-distance physics in the $B^{+}\to K^{+}\nu\bar{\nu}$ amplitude, and of new light, elusive particles produced via $B^{+}\to K^{+}a$—the two probes working independently to an excellent approximation.
+We introduce a model-independent framework to reinterpret Belle II results using only public data, analytically reconstructing the mapping between true and reconstructed kinematic variables within the statistically dominant Inclusive Tagging Analysis.
+This enables rare-decay measurements to probe light invisible particles—such as the QCD axion or axion-like particles, collectively denoted aa—without relying on internal simulations.
+Applying the method to $B^{+}\!\to\!K^{+}\nu\bar{\nu}$ yields the strongest bound on the branching fraction for $B^{+}\!\to\!K^{+}a$, improving existing limits by about a factor of nine and constraining the axion’s fundamental flavour-changing coupling to $b$ and $s$ quarks.
+The approach establishes $B^{+}\!\to\!K^{+}\nu\bar{\nu}$ as a dual probe—simultaneously testing short-distance new physics and light invisible states, the two probes working independently to an excellent approximation—and provides a general strategy for model-independent reinterpretation of collider data.
 preprint: LAPTH-036/25
-Introduction— Spin-zero, odd-parity particles below the GeV scale—collectively known as axions or axion-like particles (ALPs)—are hypothetical states actively sought through laboratory experiments [1, 2, 3], astrophysical data [4, 5, 6], cosmological probes [7], and theoretical studies [8, 9, 10].
-A central focus of this research domain is the “invisible” [11, 12, 13, 14] QCD axion [15, 16, 17, 18], henceforth the axion, which elegantly solves the strong-$CP$ problem—the unexplained suppression of a $CP$-violating QCD term.
-Given the extreme suppression, $O(10^{-10})$, neither coincidence nor anthropic reasoning provides a satisfactory explanation [19, 20].
-The axion also offers a viable Dark Matter candidate [21, 22, 23].
-Unlike generic ALPs, its mass $m_{a}$ and decay constant $f_{a}$ are linked: smaller $m_{a}$ implies weaker couplings, scaling as $1/f_{a}$.
-Existing bounds indicate that $m_{a}$ must be several orders of magnitude below the pion mass [9].
-Yet this small mass also opens the possibility that axions could be copiously produced at high-intensity colliders, without having to resort to high energies.
-One can consider decays of $K$, $D$, and $B$ mesons, or even the baryonic counterparts, decaying to fully reconstructible final states, plus the axion, which is assumed to escape without decay nor interaction within the detector volume.
-This represents the most general and conservative scenario.
-Since the QCD axion manifests as an energy and momentum imbalance—equal in magnitude, as the axion has negligible mass with respect to the lightest physical scale of the decay—it is crucial that the remaining final-state particles be reconstructible with high precision.
-Moreover, the identity and momentum of the decaying hadron must be accurately known to better control the many backgrounds involving elusive particles that can mimic the axion.
-Belle II [24] is particularly well positioned to satisfy both of these constraints.
-It collides $e^{-}$ and $e^{+}$ beams with momenta in approximately opposite directions and magnitudes of $7$ and $4$ GeV  [^1] The spread, or bite, around these momenta is tiny and completely negligible for the purposes of this paper., respectively.
-The center-of-mass energy corresponds to the $b\bar{b}$ resonance known as $\Upsilon(4S)$, which subsequently decays with near-unit probability into charged or neutral $B\bar{B}$ meson pairs.
-Signals are then studied in the decays of one of the produced mesons, $B_{\mathrm{s}ig}$, while the other meson is often used as a tag decay—allowing the kinematics of the signal decay to be fully reconstructed and the flavour of $B_{\mathrm{s}ig}$ to be identified.
-In this paper we will consider the $B^{+}\to K^{+}a$ decay  [^2] The conjugated mode with negatively-charged mesons is assumed to be included throughout our study..
-The axion-hadron interaction responsible for the decay inherits in a calculable way from the underlying axion-quark interaction, which just below the electroweak scale can be written most generally as [27]
+Introduction— Axions and axion-like particles (ALPs) are hypothetical light, spinless, pseudoscalar states actively sought through laboratory experiments [1, 2, 3], astrophysical observations [4, 5, 6], cosmological probes [7], and theoretical studies [8, 9, 10].
+A central target is the “invisible” [11, 12, 13, 14] QCD axion [15, 16, 17, 18], henceforth the axion, which elegantly solves the strong-$CP$ problem—the unexplained suppression of a $CP$-violating QCD term.
+Given the extreme smallness of this term, $O(10^{-10})$, neither coincidence nor anthropic reasoning offers a satisfactory explanation [19, 20].
+The axion is also a viable dark-matter candidate [21, 22, 23].
+Throughout this paper we denote the QCD axion or a generic ALP as $a$.
+Unlike generic ALPs, the QCD-axion’s mass $m_{a}$ and decay constant $f_{a}$ are linked: a lighter QCD axion necessarily couples more weakly—its interactions scaling as $1/f_{a}$.
+Although this relation places the QCD axion well below the pion mass [9], it simultaneously opens the door to its production at high-intensity colliders, where large event yields can compensate for its small couplings.
+In this regime, rare decays of heavy mesons such as $K$, $D$, and BB—or their baryonic counterparts—provide a unique window: the axion would manifest as missing energy and momentum, while the hadronic system, for suitably chosen channels, remains fully reconstructible.
+This defines a clean, general, and conservative search strategy capable of probing both QCD axions and generic ALPs.
+A key experimental requirement for such searches is that the visible decay products be measured with high precision and the kinematics of the parent hadron be well constrained, enabling the identification of missing energy with minimal model dependence.
+The Belle II experiment [24] is particularly well suited in this regard: it collides $e^{-}$ and $e^{+}$ beams with momenta of approximately $7$ and $4$ GeV  [^1] The spread, or bite, around these momenta is tiny and completely negligible for the purposes of this paper., producing $b\bar{b}$ pairs almost exclusively through the $\Upsilon(4S)$ resonance.
+The latter decays with near-unit probability into charged or neutral $B\bar{B}$ meson pairs.
+Signal decays are then studied in one of the two $B$ mesons, while the other is used as a tag, allowing the kinematics of the signal decay to be reconstructed and the flavour of the signal $B$ to be identified.
+In this work we consider the decay $B^{+}\to K^{+}a$  [^2] The charge-conjugate mode is assumed to be included throughout..
+The axion–hadron interaction relevant for this decay is obtained by matching the axion–quark interaction onto the low-energy chiral effective theory.
+As a result, the axion–hadron couplings are unambiguously determined—up to the standard hadronic low-energy constants at the chosen chiral order—by the underlying axion–quark couplings.
+Just below the electroweak scale, the latter can be written in the most general form as [27]
 $\mathcal{L}_{a}\supset\;$ $\frac{\partial_{\mu}a}{2f_{a}}\sum_{i,j=1}^{3}\bar{d}_{i}\gamma^{\mu}\bigl[(k_{V})_{d_{i}d_{j}}+\gamma_{5}(k_{A})_{d_{i}d_{j}}\bigl]d_{j}+\left\{d\to u\right\}$
@@
 The process $B^{+}\to K^{+}a$, on which this paper is focused, involves the complex coupling $(k_{V})_{sb}$. The calculation of this process requires the following hadronic matrix element
-$\langle K^{-}(k)|\bar{s}b|B^{-}(p)\rangle=\frac{m_{B}^{2}-m_{K}^{2}}{m_{b}-m_{s}}f_{0}(m_{a}^{2})~.$ (2)
-Here, form-factor values and conventions are taken from Ref. [29]. The relation in Eq. (2) imply the following branching ratios (BRs) [30, 31]
+$\langle K^{-}(p-q)|\bar{s}b|B^{-}(p)\rangle=\frac{m_{B}^{2}-m_{K}^{2}}{m_{b}-m_{s}}f_{0}(q^{2})~.$ (2)
+Here, form-factor values and conventions are taken from Ref. [29]. The relation in Eq. (2) implies the following branching ratio (BR) [30, 31]
 $\mathcal{B}(B^{-}\to K^{-}a)=\frac{\tau_{B}|(k_{V})_{sb}|^{2}\,|\vec{k}|m_{B}^{2}}{32\pi f_{a}^{2}}\left(1-\frac{m_{K}^{2}}{m_{B}^{2}}\right)^{2}f_{0}^{2}(m_{a}^{2})~,$ (3)
@@
 Here and henceforth, the axion-quark couplings are understood at a scale close to $m_{B}$.
-We search for the signal Eq.
-(3) in data collected by Belle II for the $B^{+}\to K^{+}\nu\bar{\nu}$ measurements.
-Belle II employs two nearly uncorrelated datasets, analyzed using the “classic” Hadronic Tagging Analysis (HTA) method and the more recent Inclusive Tagging Analysis (ITA) method, respectively [32, 33, 34].
-The ITA method achieves much larger statistics and thus dominates the sensitivity to the measurement.
-Reliably recasting this dataset for our search thus hinges on one key piece of non-public information: the mapping between the true di-neutrino invariant mass squared, $q^{2}$, which is inaccessible within the ITA method, and its reconstructed counterpart, $q^{2}_{{\mathrm{r}ec}}$, defined below.
-We show that the $q^{2}$ distribution as a function of $q^{2}_{{\mathrm{r}ec}}$ can be reliably inferred from kinematic arguments alone—i.e., without requiring internal information from the experiment.
+We search for the signal decay $B^{+}\to K^{+}a$ in data collected by Belle II for the $B^{+}\to K^{+}\nu\bar{\nu}$ measurements.
+Belle II provides two nearly uncorrelated datasets for the $B^{+}\to K^{+}\nu\bar{\nu}$ analysis: the “classic” Hadronic Tagging Analysis (HTA) and the more recent Inclusive Tagging Analysis (ITA) [32, 33, 34].
+In the HTA, the full tag-$B$ decay chain is reconstructed, allowing the di-neutrino invariant mass $q^{2}$ to be measured directly.
+In contrast, the ITA provides far larger statistics—and therefore dominates the sensitivity—but does not reconstruct $q^{2}$; instead, it uses a proxy variable $q^{2}_{\mathrm{rec}}$ obtained as discussed below.
+Recasting the ITA sample for $B^{+}\to K^{+}a$ searches therefore requires knowledge of the $q^{2}$$q^{2}_{\mathrm{rec}}$ mapping, which is normally accessible only through internal Belle II simulation.
+A first central result of this work is to show that this mapping can be derived analytically from kinematics alone, without relying on non-public experimental inputs.
+This insight enables a model-independent reinterpretation of the ITA data for light invisible states such as the axion.
 Using this mapping and publicly available efficiencies, we demonstrate that this search yields the strongest existing bounds on the magnitudes of the fundamental QCD-axion coupling $(k_{V})_{sb}$.
-Both of the aforementioned ITA and HTA methods yield a central value in excess of the SM prediction [35, 36].
-This excess motivates the following parameterization of the BR for $B^{+}\to K^{+}\nu\bar{\nu}$—differential in the squared di-neutrino invariant mass $q^{2}$
-$d\mathcal{B}_{\nu\bar{\nu}}(\mu)/dq^{2}~\equiv~\mu_{\nu\bar{\nu}}\,d\mathcal{B}_{\nu\bar{\nu},{\mathrm{S}M}}/dq^{2}~,$ (4)
-where on the r.h.s.
-we use the SM prediction [35, 37, 38], rescaled by a signal-strength factor.
-This rescaling allows one to disentangle the presence of new effects in the $B^{+}\to K^{+}\nu\bar{\nu}$ mode from those due to the two-body decay $B^{+}\to K^{+}a$, and to test whether the bound inferred on the $(k_{V})_{sb}$ coupling is affected by possible new-physics contamination in the $B^{+}\to K^{+}\nu\bar{\nu}$ sample.
-Within the ITA, Belle II cannot directly reconstruct $q^{2}\equiv m_{\nu\bar{\nu}}^{2}=(p-k)^{2}$ and thus constructs a proxy variable, $q^{2}_{{\mathrm{r}ec}}$ [32], in terms of the known $s\equiv(p_{e^{-}}+p_{e^{+}})^{2}$ and the energy of the final-state kaon in the $B\bar{B}$ rest frame (hereafter denoted with a star).
-In this frame, the signal $B$-meson momentum is known in magnitude, $|\vec{p}_{B}^{\,*}|=0.3298~{\mathrm{G}eV}$, and is distributed isotropically.
-In turn, in the $B$ rest frame (denoted with a double star), the kaon is emitted isotropically.
-One can therefore write
+This represent our second key result.
+Both the ITA and HTA measurements show central values above the Standard-Model (SM) prediction for $B^{+}\!\to\!K^{+}\nu\bar{\nu}$ [35, 36, 37, 38].
+This motivates introducing a signal-strength parameter $\mu_{\nu\bar{\nu}}$, defined through $d\mathcal{B}_{\nu\nu}(\mu)/dq^{2}\equiv\mu_{\nu\bar{\nu}}\,d\mathcal{B}_{\nu\nu,\mathrm{SM}}/dq^{2}$.
+Allowing $\mu_{\nu\bar{\nu}}$ to float ensures that possible short-distance new-physics effects in $B\!\to\!K\nu\bar{\nu}$ are not misinterpreted as, or artificially absorbed by, the two-body decay $B\!\to\!Ka$.
+This separation is essential: it renders the analysis sensitive to both types of new phenomena and establishes $B^{+}\to K^{+}\nu\bar{\nu}$ as a dual probe—simultaneously testing short-distance modifications of the $b\!\to\!s\nu\bar{\nu}$ amplitude and the production of light invisible states $a$.
+As we will demonstrate, the two effects operate independently to an excellent approximation and this represents our third key result.
+To implement this reconstruction explicitly, we recall that within the ITA, Belle II cannot directly reconstruct $q^{2}\equiv m_{\nu\bar{\nu}}^{2}=(p-k)^{2}$ and thus constructs a proxy variable, $q^{2}_{{\mathrm{r}ec}}$ [32], in terms of the energy of the final-state kaon in the $B\bar{B}$ rest frame (denoted by a star) and of the known $s\equiv(p_{e^{-}}+p_{e^{+}})^{2}=4\,E_{B}^{*2}$.
+In this frame, the signal $B$-meson momentum has a fixed magnitude, $|\vec{p}_{B}^{\,*}|=0.3298~{\mathrm{G}eV}$, and is distributed isotropically.
+In the $B$ rest frame (double star), the kaon from the two-body decay $B\to Ka$ would likewise be monochromatic for fixed $q^{2}=m_{a}^{2}$; however, its observed momentum depends on the relative orientation between the two frames, as discussed next.
+One can write
 $\begin{split}q^{2}_{{\mathrm{r}ec}}~\equiv~&\frac{s}{4}+m_{K}^{2}-\sqrt{s}E_{K}^{*}~=~q^{2}+E_{B}^{*2}-m_{B}^{2}-2\,\vec{p}_{B}^{\,*}\cdot\vec{p}_{K}^{\,*}(q^{2})\\ ~=~&q^{2}+E_{B}^{*2}-m_{B}^{2}-2|\vec{p}_{B}^{\,*}|\gamma^{**}\left(\beta^{**}E_{K}^{**}(q^{2})+|\vec{p}_{K}^{\,**}(q^{2})|\cos\theta^{**}\right)~,\end{split}$ (4)
@@
 (4) with $c^{**}\equiv\cos\theta^{**}\in[-1,1]$ uniformly distributed.
-The sub-percent-level uncertainties [39] smearing the energies and momenta that appear in Eq.
-(5) have a completely negligible impact on the resulting $q^{2}(q^{2}_{{\mathrm{r}ec}})$ distribution.
+Eq.
+(4) makes explicit the aforementioned dependence on the relative orientation between the and ∗∗ frames, showing that its lack of knowledge is structural in the ITA: while the kaon produced in $B\to Ka$ is monochromatic in the $B$-meson rest frame, reconstructing this energy ($E_{K}^{*}$ in the $q^{2}_{{\mathrm{r}ec}}$ defining equation) requires knowledge of the full decay kinematics.
+This includes the angle $\theta^{**}$, which in the ITA is not accessible.
+The monochromatic kaon energy is thus mapped onto a broadened distribution in $q^{2}_{\mathrm{r}ec}$, whose shape is fixed almost entirely by kinematics: the sub-percent-level uncertainties [39] smearing the energies and momenta that appear in Eq.
+(4) have a completely negligible impact on the $q^{2}(q^{2}_{{\mathrm{r}ec}})$ distribution.
 The physical branching ratios and the observed yields are then related in the ITA as
@@
 (4).
-In the last equality we make contact with the smearing function $f_{q^{2}_{{\mathrm{r}ec}}}$ [41, 31, 42], which is typically assumed to require Belle II simulation.
-The $q^{2}$ vs. $q^{2}_{{\mathrm{r}ec}}$ correlation implied by Eq. (6) is shown in Fig. 1.
+In the last equality of Eq.
+(5) we make contact with the smearing function $f_{q^{2}_{{\mathrm{r}ec}}}$ [41, 31, 42], which is typically assumed to require Belle II simulation.
+As discussed, this function is instead simply fixed by Eq.
+(4), shown in Fig.
+1.
 Figure 1: Correlation between the true ($q^{2}$) and the reconstructed ($q^{2}_{{\mathrm{r}ec}}$) di-neutrino invariant mass squared. The plot is normalised to $10^{5}$ observed events.
-This correlation plot may be compared with figure 5.1 of Ref.
+This correlation plot may be compared with Fig.
+5.1 of Ref.
 [43], which is the output of internal Belle II simulations.
-The two figures display good agreement, aside from an inconsequential low-$q^{2}$ tail visible in figure 5.1 of Ref.
-[43]  [^5] This tail originates from radiative decays with undetected photons.
-Strictly speaking, $B\to K\nu\bar{\nu}(+n\gamma)$ is a distinct process from $B\to K\nu\bar{\nu}$ and should be subtracted if sizeable, as in $B_{s}\to\mu^{+}\mu^{-}$[50].
-Here, however, the effect is negligible (well below 1% in Ref.
-[43]), since the lightest emitter is the $K^{+}$, which makes the exponent $b$ in ${\mathcal{B}}(B\to K\nu\bar{\nu}(+n\gamma))=\Bigl(\tfrac{2\Delta E}{m_{B}}\Bigr)^{\tfrac{\alpha}{4\pi}b}{\mathcal{B}}(B\to K\nu\bar{\nu})$ parametrically small.
-Here $\Delta E$ denotes the minimal photon energy in the $B$ rest frame required for detectability—a “theoretical” scale usually proxied by the $\sqrt{q^{2}}$ resolution.
-.
-To make the origin of the $q^{2}_{{\mathrm{r}ec}}$ smearing explicit, we recall that the relation between $q^{2}$ and $q^{2}_{{\mathrm{r}ec}}$, Eq.
-(5), introduces a uniformly distributed variable from an unmeasured angle.
-This kinematic unknown constitutes an irreducible source of uncertainty, which is not immediately transparent in simulation output.
-The good agreement between Fig.
-1 and figure 5.1 of Ref.
-[43] shows that this mechanism alone accounts for the observed correlation, with any additional effects being negligible.
-Within the HTA, instead, Belle II can directly reconstruct the $q^{2}$ of the events, so the physical BRs and observed yields are simply related as
+The two figures display good agreement, aside from an inconsequential low-$q^{2}$ tail visible in Ref.
+[43]’s figure.
+This tail originates from radiative decays with undetected photons.
+Strictly speaking, $B\to K\nu\bar{\nu}(+n\gamma)$ constitutes a distinct physical process from the non-radiative mode and would need to be subtracted if its contribution were sizeable, as is the case for $B_{s}\to\mu^{+}\mu^{-}$ [44].
+In the present case, however, this effect is negligible—well below the percent level in Ref.
+[43]  [^5] The suppression arises because the lightest charged emitter is the $K^{+}$, which makes the exponent $b$ in ${\mathcal{B}}(B\to K\nu\bar{\nu}(+n\gamma))=\Bigl(\tfrac{2\Delta E}{m_{B}}\Bigr)^{\tfrac{\alpha}{4\pi}b}{\mathcal{B}}(B\to K\nu\bar{\nu})$ parametrically small.
+Here $\Delta E$ denotes the minimal photon energy in the $B$ rest frame required for detectability—a “theoretical” scale usually proxied by the $\sqrt{q^{2}}$ resolution..
+Inclusion of the radiative tail in the referenced figure is therefore immaterial to the above comparison.
+Eq.
+(4) and the ensuing discussion put to work an important distinction between the SM decay $B^{+}\to K^{+}\nu\bar{\nu}$ and the exotic two-body decay $B^{+}\to K^{+}a$: the latter corresponds to a fixed invariant mass $q^{2}=m_{a}^{2}$, whereas the former exhibits a continuous $q^{2}$ spectrum.
+This difference is fully exploited in our analysis: it is encoded at the level of the differential branching fractions, which enter as a smooth distribution in $q^{2}$ for the di-neutrino mode and as a Dirac delta function for the axion signal.
+The apparent loss of monochromatic information in the ITA does not reflect a limitation of principle, but arises solely from the fact that the experimentally accessible variable is $q^{2}_{\mathrm{r}ec}$ rather than the true $q^{2}$.
+By contrast, within the HTA, to which we turn next, Belle II can reconstruct the full event kinematics, allowing direct access to $q^{2}$.
+In this case, the axion signal appears as a narrow peak in the $q^{2}$ spectrum, equivalent to a monochromatic kaon energy in the $B$-meson rest frame, and the analysis reduces to a standard “bump search”.
+The physical BRs and observed yields are simply related as
 $\frac{dN_{X}}{dq^{2}}~=~N_{B}\,\epsilon_{\text{HTA}}(q^{2})\,\frac{d\mathcal{B}_{X}}{dq^{2}}~,$ (6)
 where $\epsilon_{\text{HTA}}(q^{2})$ is the HTA-specific selection efficiency as a function of $q^{2}$, also available in Ref. [32].
+We conclude by emphasizing that, within our approach, the two tagging strategies exploit the same underlying kinematic information.
+The observed differences arise solely from degrees of freedom that are not experimentally accessible in the ITA, rather than from any discarded physical information.
+In particular, the distinction between two-body (signal) and three-body ($B\to K\nu\bar{\nu}$) kinematics is fully incorporated in both tagging strategies.
 Analysis— To probe the $B^{+}\to K^{+}a$ decay through the $B^{+}\to K^{+}\nu\bar{\nu}$ measurements at Belle II, we construct a likelihood ratio test statistic based solely on the public data of Fig.
@@
 Here $n^{i}_{\mathrm{o}bs}$ is the number of observed events in bin $i$, defined by $q^{2}\in[q^{2}_{i},q^{2}_{i+1})$, and $\lambda^{i}$ is the corresponding expected number of events, which depends on $\mathcal{B}(B^{+}\to K^{+}a)\equiv\mathcal{B}_{a}$ and on the signal strength $\mu_{\nu\bar{\nu}}$—the two key quantities to be determined:
-$\begin{split}\lambda^{i}(\mathcal{B}_{a},\mu_{\nu\bar{\nu}})~=~\mathcal{B}_{a}\,n^{i}_{a}+\mu_{\nu\bar{\nu}}\,n^{i}_{\nu\bar{\nu}}+(1+\sigma^{\text{bkg}}\,\theta)\,n^{i}_{\text{bkg}}~.\end{split}$ (8)
+$\begin{split}\lambda^{i}(\mathcal{B}_{a},\mu_{\nu\bar{\nu}})~=~\mathcal{B}_{a}\,n^{i}_{a}+\mu_{\nu\bar{\nu}}\,n^{i}_{\nu\bar{\nu}}+n^{i}_{\text{bkg}}~.\end{split}$ (7)
 This expression applies to either the ITA or HTA dataset, where $n^{i}_{a}$, $n^{i}_{\nu\bar{\nu}}$, and $n^{i}_{\mathrm{b}kg}$ are the yields in the $i^{\mathrm{t}h}$ bin for the signal, the $B^{+}\to K^{+}\nu\bar{\nu}$ decay, and the background, respectively.
@@
 $n^{i}_{\nu\bar{\nu},\text{ITA}}=$ $N_{B}\int\limits_{\text{$i^{\text{th}}$ bin}}dq^{2}_{{\mathrm{r}ec}}\int dq^{2}\,f_{q^{2}_{{\mathrm{r}ec}}}(q^{2})\,\epsilon_{\text{ITA}}(q^{2})\,\frac{d\mathcal{B}_{\nu\bar{\nu},{\mathrm{S}M}}}{dq^{2}}~,$ (9)
-$n^{i}_{a,\text{HTA}}=$ $N_{B}\int\limits_{\text{$i^{\text{th}}$ bin}}dq^{2}\,\epsilon_{\text{HTA}}(q^{2})\,g_{a}(q^{2})~,$ (11)
+$n^{i}_{a,\text{HTA}}=$ $N_{B}\int\limits_{\text{$i^{\text{th}}$ bin}}dq^{2}\,\epsilon_{\text{HTA}}(m_{a}^{2})\,g_{a}(m_{a}^{2})~,$ (10)
 $n^{i}_{\nu\bar{\nu},\text{HTA}}=$ $N_{B}\int\limits_{\text{$i^{\text{th}}$ bin}}dq^{2}\,\epsilon_{\text{HTA}}(q^{2})\,\frac{d\mathcal{B}_{\nu\bar{\nu},{\mathrm{S}M}}}{dq^{2}}~.$ (11)
-These expressions can be understood in terms of Eqs.
-(6)-(7), bearing in mind that for the signal case the $q^{2}$ dependence of the differential BR is a delta function with support at $q^{2}=m_{a}^{2}$, and the overall normalization $\mathcal{B}_{a}$ is factored out in Eq.
+These expressions can be understood directly from Eqs.
+(5)-(6), bearing in mind that, for the signal, the differential branching ratio is proportional to a Dirac delta function with support at $q^{2}=m_{a}^{2}$, and that the overall normalization $\mathcal{B}_{a}$ is factored out in Eq.
 (7).
-In the HTA case, one has to take into account that the invariant mass resolution of the experiment is finite.
-We do so by replacing $\delta(q^{2}-m_{a}^{2})$ with a normal distribution $g_{a}(q^{2})$ centered at the axion mass value with uncertainty taken to be $\sim(450~{\mathrm{M}eV})^{2}$, comparable to the beam-constrained mass resolution of $B$ mesons reconstructed using the hadronic tag [45]  [^6] In principle, the same smearing should also be applied to the di-neutrino channel, Eq.
+In addition, for the HTA signal contribution one must account for the finite invariant-mass resolution of the experiment.
+We implement this by replacing $\delta(q^{2}-m_{a}^{2})$ with a normal distribution $g_{a}(q^{2})$ centered at $m_{a}^{2}$, with variance taken as $\sim(450~{\mathrm{M}eV})^{2}$, comparable to the beam-constrained mass resolution of $B$ mesons reconstructed with the hadronic tag [46].
+This prescription also allows us to comment more generally on detector-induced smearing effects.
+In principle, the same smearing procedure should be applied to the di-neutrino contribution in the HTA, Eq.
 (11).
-In the axion case this treatment is essential, as it broadens the Dirac-delta shape of the $q^{2}$ distribution.
-For the di-neutrino mode, however, the impact is much milder because its $q^{2}$ spectrum is already smooth.
-Such an effect would become relevant only once a measurement of the differential $\mathcal{B}_{\nu\bar{\nu}}$ reaches comparable precision.
-In practice, the smeared and unsmeared $q^{2}$ distributions differ only marginally: a few events with small positive $q^{2}$ migrate to slightly negative reconstructed values, thus shifting between adjacent bins of the Belle II spectrum.
-Numerically, the neutrino likelihood at $\mathcal{B}_{a}=0$ remains nearly unchanged, and the bound on the axion is identical within the two-digit precision considered..
-As regards the last component of Eq.
-(8), the numbers of expected background events for each bin, $n^{i}_{\mathrm{b}kg}$, can be read out from the plots of the collaboration.
-While the uncertainty on the Montecarlo generated background events are not explicitly provided, we infer that the uncertainty on the total normalisation is respectively $\sigma_{\text{ITA}}^{\text{bkg}}\approx 1\%$ and $\sigma_{\text{HTA}}^{\text{bkg}}\approx 10\%$.
+While in the axion case this treatment is essential, in the di-neutrino case the impact is much milder because the $q^{2}$ spectrum of $B\to K\nu\bar{\nu}$ is already smooth.
+Numerically, the smeared and unsmeared distributions differ only marginally  [^6] The dominant effect is that a few events with small positive $q^{2}$ migrate to slightly negative reconstructed values, thus shifting between adjacent bins of the Belle II spectrum.
+and the resulting likelihood and axion limits are unchanged within the two-digit precision of our results (see Table 1).
+In the ITA, the situation is fundamentally different.
+The axion signal is already broadened at the kinematic level: the monochromatic kaon energy in the $B$ rest frame (or, equivalently, the Dirac-delta structure of the axion $q^{2}$ distribution) is mapped onto a smooth distribution in the experimentally accessible variable $q^{2}_{{\mathrm{r}ec}}$.
+The finite experimental resolution on $E_{K}^{*}$ therefore introduces only a subleading additional smearing, whose effect we have verified to be numerically negligible in practice.
+We next address the role of uncertainties on the selection-efficiency factors $\epsilon$ appearing in Eqs.
+(8)-(11).
+The statistical uncertainties on the efficiencies are at the percent or sub-percent level for both channels, as reported in Ref.
+[32].
+The systematic components are, in turn, approximately 3% in ITA and 16% in HTA (see Tables I and II of Ref.
+[32]).
+Since the expected signal yields depend linearly on the efficiencies, these uncertainties propagate directly to the extracted BR limits quoted in Table 1.
+Because the bounds are largely driven by the ITA sample, the corresponding variation in the ITA and combined results would affect only the third significant digit, whereas the HTA-only limits would be impacted already at the second digit.
+This is reflected in the precision with which we report the results for the two tagging strategies.
+In practice, varying the efficiencies within their quoted uncertainties does not modify the bounds at the level of precision displayed in Table 1.
+We now pause on the treatment of systematic uncertainties affecting the last component of Eq.
+(7), namely the background normalization.
+In the Belle II analysis, the impact of a given systematic uncertainty is quantified by the reduction in the total uncertainty when that source is removed from the fit, as shown in Tables I and II of Ref.
+[32].
+While the individual components cannot be summed in quadrature to reconstruct the total uncertainty, their relative magnitudes provide a useful estimate of their importance.
+From this comparison, the combined effect of the systematic uncertainties other than the background normalization is at the level of $O(50\%)$ or less relative to the background normalization contribution.
+To model this component, we reason as follows.
+Although the uncertainties on the Monte Carlo generated background events are not explicitly provided, we infer that the relative uncertainty on the total normalisation is respectively $\sigma_{\text{ITA}}^{{\mathrm{b}kg}}\approx 1\%$ and $\sigma_{\text{HTA}}^{{\mathrm{b}kg}}\approx 10\%$.
 These estimates have been obtained from the $q^{2}$ distribution and its corresponding pull shown in Figs.
@@
 [32].
-One can in fact identify the pull with $|n^{i}_{\mathrm{d}at}-n^{i}_{\mathrm{h}st}|/(\sqrt{(\sigma^{i}_{\mathrm{d}at})^{2}+(\sigma^{i}_{\mathrm{h}st})^{2}})$, where the subscripts ‘dat’ and ‘hst’ denote respectively the data and the value of the stacked histograms, and $\sigma^{i}_{\mathrm{d}at}$ the uncertainty in the data point.
-Inverting the pull distribution in terms of $\sigma^{i}_{\mathrm{h}st}$ provides an estimate of the above uncertainty.
-Thus, we introduced two nuisance parameters, i.e.
-$\theta_{\text{ITA}}$ and $\theta_{\text{HTA}}$, both constrained by a normal Gaussian distribution in the likelihood as
-${\mathcal{L}}_{2}=\frac{1}{\sqrt{2\pi}}e^{-\theta^{2}_{\text{ITA}}/2}\times\frac{1}{\sqrt{2\pi}}e^{-\theta^{2}_{\text{HTA}}/2}\,.$ (13)
-As a non-trivial check of our procedure, we derive the profile likelihood ratio as a function of the signal strength $\mu_{\nu\bar{\nu}}$ in the case the axion is absent, i.e.
-${\mathcal{B}}_{a}=0$, obtaining very consistent results as compared with the collaboration’s counterparts.
-Details are relegated to the text around Fig.
-3 in the Supplemental Sec.
-A.
+One can identify the pull as $|n^{i}_{{\mathrm{o}bs}}-n^{i}_{{\mathrm{h}st}}|/\sqrt{n^{i}_{{\mathrm{o}bs}}+(\sigma^{i}_{{\mathrm{h}st}})^{2}}$, where the subscripts ‘obs’ and ‘hst’ denote respectively the data and the value of the stacked histograms, and $\sigma^{i}_{{\mathrm{o}bs}}$ the uncertainty in the data point.
+Inverting the pull distribution in terms of $\sigma^{i}_{{\mathrm{h}st}}$, we find that the relative uncertainties $\sigma^{i}_{{\mathrm{h}st}}/n^{i}_{{\mathrm{h}st}}$ are approximately constant across bins.
+We therefore treat them as a single relative uncertainty $\sigma^{{\mathrm{b}kg}}$ on the overall background normalization, modeled as a nuisance parameter $\theta$ in $n^{i}_{{\mathrm{b}kg}}=(1+\sigma^{{\mathrm{b}kg}}\,\theta)\,\langle n^{i}_{{\mathrm{b}kg}}\rangle$.
+This can be constrained, independently in the ITA and HTA analyses, by the Gaussian likelihood terms
+${\mathcal{L}}_{2}=\frac{1}{\sqrt{2\pi}}e^{-\theta^{2}_{{\mathrm{I}TA}}/2}\times\frac{1}{\sqrt{2\pi}}e^{-\theta^{2}_{{\mathrm{H}TA}}/2}~,$ (12)
+where the expected value $\langle n^{i}_{\mathrm{b}kg}\rangle$ of the number of background events for each bin, $n^{i}_{\mathrm{b}kg}$, can be read out from Ref. [32].
+A strong test that the above procedure captures the bulk of the systematic uncertainties can be performed by studying our profile likelihood $\mathcal{L}$ as a function of the signal strength $\mu_{\nu\bar{\nu}}$ in the case where the axion is absent, i.e.
+where ${\mathcal{B}}_{a}=0$, thus making contact with the counterpart in Ref.
+[32].
+The result is shown in Fig.
+4 in Appendix A, whose two panels show the case of neglecting vs.
+including the systematic uncertainty on the background normalization.
+The figure shows that we obtain results well consistent with the collaboration’s counterparts only when the background normalization uncertainty is included.
+This indicates that our procedure captures the leading systematic effect.
+As a further test that the remaining systematics is negligible, we also calculated the same profile likelihood, with uncertainty inflated by a factor 1.5, as if we were missing a residual $O(50\%)$ component.
+The resulting profile is essentially identical to the case with non-inflated uncertainty.
 […]
@@
 $\mu_{\nu\bar{\nu}}=1$ $n_{\text{dof}}=2$ $n_{\text{dof}}=2$ $n_{\text{dof}}=3$
-$\mathcal{B}_{a}<1.1\times 10^{-6}$ $\mathcal{B}_{a}<5.1\times 10^{-6}$ $\mathcal{B}_{a}<1.4\times 10^{-6}$
-$|(F_{V})_{sb}|>8.4\times 10^{8}$ GeV $|(F_{V})_{sb}|>3.9\times 10^{8}$ GeV $|(F_{V})_{sb}|>7.6\times 10^{8}$ GeV
+$\mathcal{B}_{a}<1.1\times 10^{-6}$ $\mathcal{B}_{a}<5\times 10^{-6}$ $\mathcal{B}_{a}<1.4\times 10^{-6}$
+    $|(F_{V})_{sb}|>8.4\times 10^{8}$ GeV     $|(F_{V})_{sb}|>4\times 10^{8}$ GeV     $|(F_{V})_{sb}|>7.6\times 10^{8}$ GeV
 $\mu_{\nu\bar{\nu}}\neq 1$ $n_{\text{dof}}=3$ $n_{\text{dof}}=3$ $n_{\text{dof}}=4$
-$\mathcal{B}_{a}<1.0\times 10^{-6}$ $\mathcal{B}_{a}<6.2\times 10^{-6}$ $\mathcal{B}_{a}<1.2\times 10^{-6}$
-$|(F_{V})_{sb}|>8.9\times 10^{8}$ GeV $|(F_{V})_{sb}|>3.6\times 10^{8}$ GeV $|(F_{V})_{sb}|>8.0\times 10^{8}$ GeV
-Table 1: Upper limits on the axion BR and lower limits on the coupling-rescaled Peccei-Quinn scale at 90% C.L. for the different cases considered. All limits assume $m_{a}=0$.
-The different cases described and the resulting bounds are summarized in Table 1, which assumes $m_{a}=0$.
+$\mathcal{B}_{a}<1.0\times 10^{-6}$ $\mathcal{B}_{a}<6\times 10^{-6}$ $\mathcal{B}_{a}<1.2\times 10^{-6}$
+$|(F_{V})_{sb}|>8.9\times 10^{8}$ GeV $|(F_{V})_{sb}|>4\times 10^{8}$ GeV $|(F_{V})_{sb}|>8.0\times 10^{8}$ GeV
+Table 1: Upper limits on the axion BR and lower limits on the coupling-rescaled Peccei-Quinn scale at 90% C.L.
+for the different cases considered.
+All limits assume $m_{a}=0$.
+See text for more details on the significant digits quoted for the different tagging strategies.
+The different cases described and the resulting bounds are summarized in Table 1, which assumes $m_{a}=0$  [^7] Ref.
+[53] obtains a bound slightly above $1\times 10^{-6}$ for the branching ratio in the low-$m_{a}$ regime, based on a projected dataset of 0.5 ab-1.
+Since Ref.
+[53] predates Ref.
+[32], a detailed numerical comparison of the limits is not straightforward.
+We note however that we achieve a very similar sensitivity—see results in Table 1—using 0.362ab-1 of actual data, dominated by the ITA selection..
 Results for a general value of the axion mass are shown in Fig.
@@
 First, as one would expect, the bound is weakened by an order of magnitude in the $m_{\nu\bar{\nu}}$ bins where data and the SM $B^{+}\to K^{+}\nu\bar{\nu}$ prediction are in tension.
-These discrepant bins can actually be used to set both an upper and a lower bound on $\mathcal{B}_{a}$ [47, 31, 42], as shown in Fig.
-4 in the Supplemental Sec.
-A.
+These discrepant bins can actually be used to set both an upper and a lower bound on $\mathcal{B}_{a}$ [49, 31, 42, 50], as shown in Fig.
+5 in Appendix A.
 Second, and less obviously, the bound depends only marginally on the assumption of new physics in $\mathcal{B}_{\nu\bar{\nu}}$—namely, on whether $\mu_{\nu\bar{\nu}}$ is set to unity or left floating.
@@
 As shown in Fig.
-5 of the Supplemental Sec.
-A, the resulting bound has only a mild dependence on the specific value of the di-neutrino signal strength—even for huge beyond-SM contributions to the $B^{+}\to K^{+}\nu\bar{\nu}$ amplitude.
+6 of Appendix A, the resulting bound has only a mild dependence on the specific value of the di-neutrino signal strength—even for huge beyond-SM contributions to the $B^{+}\to K^{+}\nu\bar{\nu}$ amplitude.
 Figure 2: Upper limit on $\mathcal{B}_{a}$ (upper row) and lower limit on the coupling-rescaled Peccei-Quinn scale $(F_{V})_{sb}$ (lower row) at 90% CL as a function of the mass $m_{a}$, and with $\mu_{\nu\bar{\nu}}$ set to unity (blue line) or left floating (red).
+Figure 3: QCD axion mass $m_{a}$ as a function of the effective flavor-violating coupling $(k_{V})_{sb}$, shown for two representative lower bounds on $(F_{V})_{sb}$ derived in this work.
+The QCD axion lies effectively at $m_{a}\simeq 0$ on collider scales, so the resulting limits correspond to vertical constraints in axion mass vs.
+coupling parameter space.
 The BR limit in Fig.
@@
 As the table shows, we obtain $|(F_{V})_{sb}|\geq 0.9\times 10^{9}$ GeV for the QCD axion, which improves by one order of magnitude the pre-existing bound [52].
-In conclusion, we reinterpret the $B^{+}\to K^{+}\nu\bar{\nu}$ measurement as a probe of $B^{+}\to K^{+}a$, with $a$ an axion ($m_{a}\simeq 0$) or ALP ($m_{a}>0$).
-We derive analytically the mapping between the measured $q^{2}_{\text{rec}}$ and the true $q^{2}$, allowing the Inclusive Tag Analysis to constrain axion emission without relying on internal simulations.
-Using public efficiencies, we set the strongest existing limit on the Peccei–Quinn scale rescaled by the fundamental axion coupling relevant to our channel.
-The result is largely independent of the assumed $\mu_{\nu\bar{\nu}}$, making $B^{+}\to K^{+}\nu\bar{\nu}$ a dual probe of short-distance new physics in the $B^{+}\to K^{+}\nu\bar{\nu}$ amplitude and of the production of an undetected axion.
-This approach can be generalized to $B^{+}\to K^{*+}\nu\bar{\nu}$—and to yet other channels—as soon as Belle II measurements become available.
+We now close the circle with the discussion in the Introduction concerning the distinction between the theoretically motivated but experimentally challenging QCD axion, and generic ALPs, whose appeal lies mostly in their experimental accessibility.
+With our approach (based on a fully reproducible analysis strategy) we obtain novel bounds on the QCD axion—despite the fact that the QCD axion is generally much more challenging than ALPs at colliders, since its mass scale lies far below any collider “scales”, as clarified below.
+More quantitatively, the axion mass is related to the $(F_{V})_{sb}$ scale as
+$m_{a}=\frac{\sqrt{2}F_{\pi}m_{\pi}}{(F_{V})_{sb}(k_{V})_{sb}}\frac{\sqrt{m_{u}m_{d}}}{m_{u}+m_{d}}~,$ (13)
+where $F_{\pi}=93\sqrt{2}~{\mathrm{M}eV}$.
+Our analysis yields lower bounds on $(F_{V})_{sb}$, with the weakest at about $4\times 10^{8}$ GeV and the strongest close to $9\times 10^{8}$ GeV, depending on the treatment of the di-neutrino signal strength and on the tagging strategy (hadronic or inclusive) (see Table 1).
+One can therefore study the corresponding range of $m_{a}$ values as a function of $|(k_{V})_{sb}|\lesssim 10$, as imposed by perturbativity.
+The result is shown in Fig.
+3.
+The figure shows that, over many orders of magnitude in the axion mass, the QCD axion is effectively located at $m_{a}\simeq 0$ on collider scales—meaning, collectively, the physical effective scales of the processes studied as well as the detection-induced resolution scales.
+Consequently, our limits act as essentially vertical constraints in axion parameter spaces of mass vs.
+coupling, constraining the effective flavor-violating coupling $(k_{V})_{sb}$ independently of $m_{a}$.
+This representation clarifies how our results should be interpreted within the broader axion-search landscape, which includes experiments that are sensitive to the axion mass range in our Fig.
+3: while collider experiments are insensitive to the tiny QCD axion mass itself, they can nevertheless impose meaningful and novel constraints on its flavor structure.
+In conclusion, we reinterpret the $B^{+}\to K^{+}\nu\bar{\nu}$ measurement as a probe of the two-body decay $B^{+}\to K^{+}a$, with $a$ an axion ($m_{a}\simeq 0$) or a generic ALP ($m_{a}>0$).
+A central ingredient of this work is the analytic reconstruction of the mapping between the measured $q^{2}_{\mathrm{rec}}$ and the true $q^{2}$, allowing the Inclusive Tag Analysis to be fully used without relying on internal simulations.
+Combined with publicly available efficiencies, this yields the strongest existing limit on the Peccei–Quinn scale rescaled by the fundamental vector flavour-changing coupling relevant to our channel.
+A key outcome is that this bound is largely insensitive to assumptions about the di-neutrino signal strength $\mu_{\nu\bar{\nu}}$.
+As a result, the decay $B^{+}\to K^{+}\nu\bar{\nu}$ acts as a dual probe: it simultaneously constrains short-distance new physics in the $b\!\to\!s\nu\bar{\nu}$ amplitude and the production of light invisible states such as the axion—with the two effects remaining statistically independent to excellent accuracy.
+The strategy developed here can be applied directly to $B^{+}\to K^{*+}\nu\bar{\nu}$ and, more generally, to other Belle II channels as soon as the corresponding measurements become available, providing a model-independent framework for public-data reinterpretation in rare decays.
 Acknowledgments— We thank Joel C.
@@
 This work has received funding from the French ANR, under contracts ANR-19-CE31-0016 (‘GammaRare’) and ANR-23-CE31-0018 (‘InvISYble’), that we gratefully acknowledge.
-A Supplementary plots
+Appendix A Supplementary plots
 In this part we collect additional plots which are not necessary to understand the results in the main text, but further corroborate them, or can be used to perform comparisons.
@@
 In Fig.
-5, we evaluate the bound on $\mathcal{B}_{a}$ for fixed values of $\mu_{\nu\bar{\nu}}$ different from unity, illustrating the mild dependence of the resulting bound on the assumed value of the di-neutrino signal strength.
-Figure 3: Twice the negative profile log-likelihood ratio as a function of $\mu_{\nu\bar{\nu}}$ for the ITA (red), HTA (blue) or combined (black) analyses.
-The solid line refers to our results while the dashed line are the results from Belle II in Fig.
+6, we evaluate the bound on $\mathcal{B}_{a}$ for fixed values of $\mu_{\nu\bar{\nu}}$ different from unity, illustrating the mild dependence of the resulting bound on the assumed value of the di-neutrino signal strength.
+Figure 4: Twice the negative profile log-likelihood ratio as a function of $\mu_{\nu\bar{\nu}}$ for the ITA (red), HTA (blue), and combined (black) analyses.
+Solid lines correspond to our results, while dashed lines reproduce the Belle II results shown in Fig.
 16 of Ref.
 [32].
-The grey dashed lines refer to $1,2,3\sigma$ $\chi^{2}$ values with 1 degree of freedom.
+The horizontal grey dashed lines indicate the $1,\,2,\,3\sigma\chi^{2}$ thresholds for one degree of freedom.
+(Left) The number of background events in each bin is fixed to its expected value $\langle n^{i}_{{\mathrm{b}kg}}\rangle$.
+(Right) The background normalization is floated as described in the text.
 Figure 5: Regions $\mathcal{B}_{a}$ at 68% (green), 95% (yellow) and 99% (orange) CL and best fit value (solid green line) of $\mathcal{B}_{a}$ as a function of the mass with $\mu_{\nu\bar{\nu}}$ fixed to 1 (upper row) or left floating (lower row).